when electron in hydrogen atom jumps frm second orbit to first orbit, the wavelength of radiation emitted is X. when electron jumps frm third orbit to first orbit , the wavelength of emitted radiation will be......
a) (27/32)X
b) (32/27)X
c) (2/3)X
d) (3/2)X

Grade:12

1 Answers

Sunil Kumar FP
10 years ago
we know that 1/X=RZ^2(1/n1^2-1/n2^2)
eq(1) 1/X=R(1-1/4)
let the wavelength be y
1/Y=R(1-1/9)
solving these two eqution
Y=27X/32

thanks and regards
sunil kr
askIItian faculty